# Analysis of a finite element method for second order uniformly elliptic PDEs in non-divergence form

### v2

@@ 4.3. Optimal convergence of numerical solution of FEM for Hamilton-Jacobi-Bellman equation
$\|u_{h}\|_{W^{2,p}_{h}(\Omega)}\leq C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}.$
@@
(94) $\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}\leq C\big(\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big),$
@@
(The condition (93) includes non-trivial example) We investigate that under which restrictions, the condition (93) can be valid for [40, Example $1$].
We recall that [40, Example $1$] is defined in two dimensional domain, such that
@@
We notice that $\epsilon\in(\kappa,1)$ where $\kappa\in[0,1)$ is not known explicitly. Then by the continuity argument, there exists $\beta\in(0,\frac{\pi}{3}]$ such that (96) with $\epsilon=\kappa$ holds for some $\lambda\in(\frac{8}{7}c_{0},2c_{0})$ and for any $\theta\in[0,\beta]$. Therefore, we can conclude that conservatively speaking, the condition (93) can be satisfied by [40, Example $1$] with probably more restricted region of the parameter $\theta$ (a non-trivial closed subinterval contained in $0,\frac{\pi}{3}]$).
@@
According to the proof of [40, Lemma $1$], we have that for any $\boldsymbol{x}\in\Omega$ almost everywhere,

### v3

@@ 4.3. Optimal convergence of numerical solution of FEM for Hamilton-Jacobi-Bellman equation
$\|u_{h}\|_{W^{1,p}(\Omega)}+\|u_{h}\|_{W^{2,p}_{h}(\Omega)}\leq C\|\sup_{\tilde{\alpha}\in\tilde{\Lambda}}|f^{\tilde{\alpha}}|\|_{L^{p}(\Omega)}.$
@@
(101) $\|u_{h}-u\|_{W^{1,p}(\Omega)}+\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}$
$\leq$ $C\big(\|u-\chi_{h}\|_{W^{1,p}(\Omega)}+\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big),$
@@
(The condition (100) includes non-trivial example) We investigate that under which restrictions, the condition (100) can be valid for [43, Example $1$].
We recall that [43, Example $1$] is defined in two dimensional domain, such that
@@
We notice that $\epsilon\in(\kappa,1)$ where $\kappa\in[0,1)$ is not known explicitly. Then by the continuity argument, there exists $\beta\in(0,\frac{\pi}{3}]$ such that (103) with $\epsilon=\kappa$ holds for some $\lambda\in(\frac{8}{7}c_{0},2c_{0})$ and for any $\theta\in[0,\beta]$. Therefore, we can conclude that conservatively speaking, the condition (100) can be satisfied by [43, Example $1$] with probably more restricted region of the parameter $\theta$ (a non-trivial closed subinterval contained in $0,\frac{\pi}{3}]$).
@@
According to the proof of [43, Lemma $1$], we have that for any $\boldsymbol{x}\in\Omega$ almost everywhere,